OpenStudy (anonymous):

Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)? A. y = x + 1 B. y = x + 1 C. y = x – 1 D. y = x – 1

OpenStudy (anonymous):

how do you solve this?

OpenStudy (anonymous):

as it is a MC question, what we can do is to sub x=-4 and y=-3. Then, check if LHS =RHS if yes, it is the answer

OpenStudy (anonymous):

Of course, you also have to check for x=-2 and y=-1

OpenStudy (anonymous):

i know the answer is B but i dont understand how to solve it and i dont understand what your saying

OpenStudy (anonymous):

Let me provide you with a detailed solution. (also suit for long questions) as it is a st. line passes through (-2, -1) and (-4, -3) slope between these two points is also the slope of the st. line. slope=(-2-(-4)) /(-1-(-3)) = 2/2=1 y-(-1) / x-(-2) =1 y+1 / x+2 =1 y+1 =x+2 y=x+1

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

you are welcome :D